2x+3y=19
X+5y=27
How do you solve problems such as this one? Thank you!

Answer by Alan3354(69443)

2x+3y=19
X+5y=27
--------
Multiply the 2nd equation by 2
--
2x+ 3y = 19
2x+10y = 54
----------------- Subtract
-7y = -35
y = 5
etc